# MMAO-Bench: MultiModal All in One Benchmark Reveals Compositional Law between Uni-modal and Omni-modal in OmniModels

### v1

@@ Abstract
Multimodal Large Languages models have been progressing from uni-modal understanding toward unifying visual, audio and language modalities, collectively termed omni models. However, the correlation between uni-modal and omni-modal remains unclear, which requires comprehensive evaluation to drive omni model’s intelligence evolution. In this work, we propose a novel, high quality and diversity omni model benchmark, MultiModal All in One Benchmark (MMAO-Bench), which effectively assesses both uni-modal and omni-modal understanding capabilities. The benchmark consists of 1880 human curated samples, across 44 task types, and a innovative multi-step open-ended question type that better assess complex reasoning tasks. Experimental result shows the compositional law between cross-modal and uni-modal performance and the omni-modal capability manifests as a bottleneck effect on weak models, while exhibiting synergistic promotion on strong models.
MMAO-Bench: MultiModal All in One Benchmark Reveals Compositional Law between Uni-modal and Omni-modal in OmniModels
@@ 1 Introduction
Multimodal artificial intelligence has undergone extensive researches in visual language model and audio language model, with current advancements progressing toward unifying visual, audio and language modalities, collectively termed omni models. The evaluation paradigm for these models has consequently expanded from assessing uni-modal understanding capabilities (i.e. visual understanding, audio understanding) to the next-level of intelligence (i.e. omni-modal understanding).
@@
Existing omni-model evaluation benchmarks remain relatively scarce and focus on different aspects. For example, some prioritize image comprehension(Li et al., 2024b), others emphasize video understanding(Hong et al., 2025), while a subset concentrates on speech interaction(Gong et al., 2024). Notably, existing datasets are exclusively English-centric, lacking evaluation benchmarks for Chinese linguistic contexts.
@@
The ideal omni-model should simultaneously preserve visual understanding capabilities (e.g., MMBench(Liu et al., 2024a)/MathVista(Lu et al., 2024)/MVBench(Li et al., 2024a)), speech interaction proficiency (e.g., MMAU(Sakshi et al., 2025)), and cross-modal integration capacity (e.g., OmniBench(Li et al., 2024b)/WorldSense(Hong et al., 2025)). Current evaluation paradigms employ disjointed benchmark suites for separate capability testing, creating resource-intensive evaluation processes and disconnected modality assessments. Beyond uni-modal, omni-modal capability introduces advanced challenges across image, video and audio modality. However, 77% questions from WorldSense are solvable without video or audio, and 25% questions from OmniBench contain erroneous answers. These designs limit the evaluation and analysis of omni models’ abilities.
Addressing theses limitations, we propose a novel, high quality and diversity omni model benchmark, MultiModal All in One Benchmark (MMAO-Bench), that enables comprehensive model assessment and pushing omni model to the next-level of intelligence.
Main Contributions:
1. Propose the first MultiModal All in One Benchmark evaluation framework specifically designed for omni models, which efficiently assesses both uni-modal and omni-modal understanding capabilities. This framework verifies the compositional law between omni-modal and uni-modal capability. For weaker models, this manifests itself as a "short-board effect" (where performance is constrained by the weakest modality), while stronger models demonstrate emergent ability across modalities.
2. Establish a systematic and unified dataset framework, featuring complete manual annotation to ensure high quality and diversity. This approach effectively prevents data contamination while better aligning with practical application scenarios. As a result, the MMAO-Bench comprises 44 different types of tasks and is 100% passing the quality check, which significantly surpassing existing datasets.
3. Beyond conventional multiple-choice formats, the evaluation incorporates innovative Multi-Step Open-Ended Question tasks to show a more realistic and discriminative evaluation result on complex reasoning.
2 Related Work
2.1 Uni-Modal Benchmarks
Based on large language models, vision language models (VLM) (Bai et al., 2025; Xiaomi, 2025; Zeng et al., 2025) and audio language models (ALM) (Ding et al., 2025; Wu et al., 2025) introduce the general intelligence to vision modality and audio modality respectively. Various uni-modal benchmarks conduct comprehensive evaluations on VLMs (Liu et al., 2024a; Lu et al., 2024; Wang et al., 2024a, b; Liu et al., 2024c; Mathew et al., 2021; Ouyang et al., 2024; Li et al., 2024a; Wu et al., 2024; Liu et al., 2024b; xAI, 2023; Xiao et al., 2021; Huang et al., 2025; Hu et al., 2025; Fu et al., 2024) and ALMs (Ardila et al., 2019; Wang et al., 2021; Yang et al., 2024; Ao et al., 2024). For image modality, MMBench(Liu et al., 2024a) proposed a systematically designed benchmark to evaluate general image understanding on 20 different tasks. Focused on mathematic, MathVision(Wang et al., 2024a) collected questions from 19 mathematic competitions to evaluate VLMs complex reasoning ability. In addition to above, OCRBench(Liu et al., 2024c) supplied the evaluation on text recognition and document understanding. For video modality, MVBench(Li et al., 2024a) aggregated 11 public video benchmarks and incorporated data enhancement process to cover 20 dynamic video understanding tasks. To complement the long video understanding area, LongVideoBench(Wu et al., 2024) introduces hourly video materials to evaluate the information retrieval ability from long context. For audio modality, MMAU(Sakshi et al., 2025) provides general audio understanding assessment across speech, sounds and music domains, featuring diverse audio samples. There’s massive uni-modal benchmarks covering diverse model abilities on vision modality and audio modality separately.
2.2 Omni-Modal Benchmarks
Omni models have arisen in recent years(Comanici et al., 2025; Xu et al., 2025b; AI et al., 2025; Li et al., 2025), as the pioneer, Gemini(Comanici et al., 2025) shows a strong ability in understanding both vision and audio, while qwen-3-omni(Xu et al., 2025b) provides leading performance in open-sourced models. However, there’s less omni-modal benchmarks that can evaluate the modality combination across image, video and audio. OmniBench(Li et al., 2024b) inserted audio as a context into the image understanding task and made up an omni-modal benchmark, while the data quality needs further improvement. WorldSence(Hong et al., 2025) emphasized audio-visual data in real world scenarios with high data quality, while most audio-visual questions can be solved by audio or video alone, which can’t assess the cross-modality ability. Other datasets focus on audio (Gong et al., 2024) or video (Zhou et al., 2025) and cover limited task types.
Due to the limited quality and coverage of existing benchmarks, we introduce a new benchmark MMAO-Bench which comprise 1250 human crafted questions for omni-modal and 630 enhanced samples for uni-modal, across 44 task types with 98% cross-modality solvable. The materials are collected from in-house dataset with manual annotation which prevents data contamination while better aligning with practical application scenarios. Beyond conventional multiple-choice formats, the evaluation incorporates innovative multi-step open-ended question tasks to show a more realistic and discriminative evaluation result on complex reasoning. Besides the self-constructed omni-modal dataset, we construct a uni-modal dataset by aggregating public uni-model datasets systematically and design a clustering-guided sampling method to achieves both evaluation efficiency and consistency. In this way, our benchmark can have a thoroughly evaluation that requires omni models to keep uni-modal abilities and obtain omni-modal abilities at the same time.
@@ 2.2 Omni-Modal Benchmarks
Daily-Omni V+A - - public 6 1197 MC EN
MMAO-Bench I+V+A I/V/A 100% 98% 90%private 44 1880 MC/MO EN/ZH
Table 1: Comparison of MultiModal Benchmarks. Where I,A,V and T indicate image, audio, video and text modality. The accuracy of question answer pairs and the percent of questions needing omni-modal to solve is shown in Acc. and Solvable. Source indicates where the materials are collected from, public datasets or private datasets that can prevent data contamination. For QA type, MC, SO and MO indicates multi-choise QA, single-open question QA and multi-open question QA. EN and ZH indicates english and chinese.
@@ 3 Method
In this section, we first introduce the comprehensive design of model ability system in Section.3.1. The the high-quality data construction pipeline are introduced in Section.3.2. The multi-step open-ended questions are introduced in Section.3.4. Finally, to improve the evaluation efficiency, we introduce a general evaluation dataset compression method in Section.3.3.2.
3.1 Model Ability System
As shown in Figure.1, the model’s capabilities are systematically categorized into two primary dimensions: Perception and Reasoning.
Perception Layer Structured through six recognition types: Object Perception, Attribute Perception, Scenario Perception, Spatial Perception, Cross- Modal Conversion, Semantic Understanding. In addition, we incorporate Cross-Modal Alignment evaluation to assess information synchronization across modalities.
Reasoning Layer extends conventional reasoning categories (General, STEM, Code) with: Spatial Reasoning, Temporal Reasoning, and Complex Reasoning (multi-conditional, multi-step problem solving).
This ability system combines uni-modal and cross-modal abilities which enables cross-modal measurement that is particularly critical for advanced omni-models. Based on this system, we create a diversity dataset with 44 task types distributed as shown in Figure.2.
Figure 1: Illustration of the omni-modal ability system proposed in MMAO-Bench.
Figure 2: Data distribution in model abilities and corresponding task types.
@@ 3.2 Omni-modal Dataset Construction
We have established a standardized and efficient data construction process (Figure.3) that produces high-quality and high-diversity dataset.
Figure 3: Data Construction Pipeline
@@
Preparation: (1) Task Design. Deconstruct cross-modal capability systems and task types. Details can be seen in Figure.2. (2) Material Collection. Our materials have the following three characteristics: a. Diverse Sources. The majority of our materials are real-world photos and videos collected through crowdsourcing, with another portion sourced from copyright-free website. Additionally, a small fraction comes from high-quality public datasets such as MMVU, LongVideoBench, and VideoVista. b. Rich and Diverse Topics. Our materials cover a broad spectrum of subjects, including society, culture, art, life, literature, science, and so on. c. Live-Recorded Audio. Apart from background sounds and music, all dialogue is recorded by human speakers. With over 20 participants in the recording process, the audio features are rich and closely reflect the diverse vocal characteristics of the real world. (3) Material filtering. Eliminate meaningless, illogical, and low-quality materials, and categorize the remaining materials by theme to create a material library. Additionally, label the materials with more detailed information such as subject, event, scene, and style to facilitate subsequent annotators to quickly find matching materials.
@@
QA Annotation: (1) Annotators. It mainly includes two parts: human experts and high-quality crowd-sourced users. Human experts have extensive experience in cross-modal data construction and annotation, a deeper understanding of model capabilities, and thus ensure higher professionalism and specificity in the data they construct. Most crowd-sourced users are college students with rich experience in multi-modal model interactions and diverse professional backgrounds, providing data with better authenticity and diversity. (2) Annotation process. First, annotators clarify the required image/video features based on task type definitions and filter appropriate materials from existing libraries using tags. Second, following data construction requirements, they then design prompts and corresponding answers. Third, to enhance data authenticity, all dialogue audio is recorded manually. Through this workflow, we ultimately generate complete QA pairs encompassing three modalities: visual, auditory, and textual.
@@
Quality Inspection: We have established a multi-stage, cyclically validated quality assurance system composed of automated tools and manual review. Each question undergoes at least three rounds of independent quality inspection to maximize data quality. First, a preliminary model check is conducted to filter out cases with ambiguous questions, non-unique answers, or those that do not conform to the task type. Secondly, through modality ablation experiments, we remove one modality of information from the QA pair to see if the model can answer based solely on the remaining information. If the question becomes unanswerable or ambiguous after removing any one modality, it proves the three-modality solvability of the data. Finally, manual quality inspection and revision are performed.
@@ 3.3.1 Quality Improvement
Existing public uni-modal datasets are bothered by data leakage issue. To verify the influence, we adopt privatization improvement on the widely used public dataset MMBench. As shown in Figure4, experiments show that the performance of models have better distinguishability after dataset improvement, reflecting the true capability differences between models. Therefore, for uni-modal data, we also follow the aforementioned construction process for self-construction datasets.
Figure 4: Comparison result of privatization improvement. After improvement, the performances among models are more progressively and distinguishable.
In addition to self-constructed data, we also selected some multimodal data from public datasets to supplement in terms of capability items and data types. (Data mainly comes from AV-Odyssey and WorldSense, accounting for 11% of the total). The specific selection logic is as follows:
1. Comprehensiveness: In terms of capabilities, focus on supplementing the perception part with a relatively low self-construction proportion, while also adding some reasoning questions; in terms of data types, prioritize selecting video + audio modality combinations with a lower self-construction proportion for supplementation, followed by image + audio.
2. Diversity: Supplement material types not covered in self-construction data to enhance diversity.
3. High Quality: Pay attention to the quality of datasets during screening (whether uni-modal answers are possible, accuracy, reasonableness, etc.). For most world senses that can be answered with uni-modal data, they will not be included.
4. Discriminative: Pay attention to the performance of this data subset on the model during screening, and remove overly difficult subsets with little discriminative power.

### v2

# UNO-Bench: A Unified Benchmark for Exploring the Compositional Law Between Uni-modal and Omni-modal in OmniModels

@@ Abstract
Multimodal Large Languages models have been progressing from uni-modal understanding toward unifying visual, audio and language modalities, collectively termed omni models. However, the correlation between uni-modal and omni-modal remains unclear, which requires comprehensive evaluation to drive omni model’s intelligence evolution. In this work, we propose a novel, high quality and UNified Omni model benchmark, UNO-Bench, which effectively assesses both UNi-modal and Omni-modal capabilities. The benchmark consists of 3730 human curated samples, with 98% cross-modality solvability, across 44 task types, and an innovative multi-step open-ended question type for assessing complex reasoning. Besides, a general scoring model supporting 6 question types is proposed for automated evaluation with 95% accuracy. Experimental result shows the Compositional Law between omni-modal and uni-modal performance and the omni-modal capability manifests as a bottleneck effect on weak models, while exhibiting synergistic promotion on strong models. Our code and data are available at GitHub.
UNO-Bench: A Unified Benchmark for Exploring the Compositional Law Between Uni-modal and Omni-modal in OmniModels
@@ 1 Introduction
Multimodal artificial intelligence has undergone extensive researches in visual language model and audio language model, with current advancements progressing toward unifying visual, audio and language modalities, collectively termed omni models. The evaluation paradigm for these models has consequently expanded from assessing uni-modal understanding capabilities (i.e. visual understanding, audio understanding) to the next-level of intelligence, omni-modal understanding.
@@
Existing omni model evaluation benchmarks remain relatively scarce and focus on different aspects. For example, some prioritize image comprehension(Li et al., 2024b), others emphasize video understanding(Hong et al., 2025), while a subset concentrates on speech interaction(Gong et al., 2024). Notably, existing datasets are exclusively English-centric, lacking evaluation benchmarks for Chinese linguistic contexts.
@@
The ideal omni model should simultaneously preserve visual understanding capabilities (e.g., MMBench(Liu et al., 2024b)/MathVista(Lu et al., 2024a)/MVBench(Li et al., 2024a)), speech interaction proficiency (e.g., MMAU(Sakshi et al., 2025)), and cross-modal integration capacity (e.g., OmniBench(Li et al., 2024b)/WorldSense(Hong et al., 2025)). Current evaluation paradigms employ disjointed benchmark suites for separate capability testing, creating resource-intensive evaluation processes and disconnected modality assessments. Beyond uni-modal, omni-modal capability introduces advanced challenges across image, video and audio modality. However, 77% questions from WorldSense are solvable without vision or audio, and 25% questions from OmniBench contain erroneous answers. These issues limit the evaluation and analysis of omni models’ capabilities.
@@
Daily-Omni V+A 94% 59% public 6 1197 MC EN
UNO-Bench-omni I+V+A -  100% 98% 90% private 44 1250 MC/MO EN/CH
UNO-Bench-uni - I/V/A 86% 97% 42% private 44 2480 MC EN/CH
Table 1: Comparison of MultiModal Benchmarks, with I, A, V, and T representing image, audio, video, and text modalities, respectively. It reports on the accuracy of question-answer pairs and the percentage of questions requiring omni-modal solutions, labeled as Acc. and Solvable. The Source category specifies the origin of the materials. Private sources, as opposed to public ones, can prevent data contamination. QA types include MC for multi-choice questions and MO for multi-step open-ended questions. EN and CH denote English and Chinese languages. UNO-Bench includes 1250 human-curated samples in the omni-modal section (referred to as -omni) and 2480 enhanced samples in the uni-modal section (referred to as -uni).
Due to the limited quality and coverage of existing benchmarks, we introduce a novel and unified benchmark UNO-Bench. As shown in Figure.1, the materials are collected from human crafting which prevents data contamination while better aligning with real-world scenarios. Beyond conventional multiple-choice questions, the evaluation adopts an innovative Multi-Step Open-Ended Question type to show a more realistic and discriminative evaluation result on complex reasoning. Besides the human crafted dataset, we incorporate existing uni-modal datasets by aggregating them systematically and design a clustering-guided sampling method to achieves both evaluation efficiency and consistency. In this way, our benchmark involves a comprehensive assessment necessitating omni models to maintain their uni-modal ability while simultaneously acquiring omni-modal capability.
Main Contributions:
1. Propose the first UNified Omni model benchmark, UNO-Bench, which efficiently assesses both UNi-modal and Omni-modal understanding capabilities. UNO-Bench verifies the Compositional Law between omni-modal and uni-modal capability. The omni-modal capability acts as a bottleneck effect on weaker models, but shows synergistic enhancement on stronger models.
2. Establish a high quality and diversity dataset construction pipeline including human-centric process and automated data compression. As as result, UNO-Bench comprises 1250 human curated samples for omni-modal with 98% cross-modality solvability, and 2480 enhanced samples for uni-modal, across 44 task types and 6 question types. The human-created novel dataset is well-suited to real-world scenarios, particularly within the Chinese context, whereas the automatically compressed dataset offers a 90% increase in speed and maintains 98% consistency across 18 public benchmarks. Its comprehensive quality and efficiency significantly surpasses existing datasets.
3. Beyond conventional multiple-choice question type, the evaluation incorporates innovative Multi-Step Open-Ended Question (MO) to show a more realistic and discriminative evaluation result on complex reasoning especially for multi-step reasoning across modalities. For automated evaluation, a General Scoring Model is proposed to support 6 kinds of question types with 95% accuracy on OOD models and benchmarks.
Figure 1: Illustration of the unified ability taxonomy proposed in UNO-Bench. (a) exemplarily presents the way to unify uni-modal and omni-modal samples. (b) shows the three-layer hierarchical structure of the unified ability taxonomy and its distribution in UNO-Bench. Among the 13 abilities, (c) demonstrates the $4$ diverse task types that cover a large range of real world scenarios.
2 Related Work
2.1 Uni-Modal Benchmarks
Based on large language models, vision language models (VLM) (Bai et al., 2025; Xiaomi, 2025; Zeng et al., 2025) and audio language models (ALM) (Ding et al., 2025; Wu et al., 2025) introduce the general intelligence to vision modality and audio modality respectively. Various uni-modal benchmarks conduct comprehensive evaluations on VLMs (Liu et al., 2024b; Lu et al., 2024a; Wang et al., 2024b, c; Liu et al., 2024d; Mathew et al., 2021; Ouyang et al., 2024; Li et al., 2024a; Wu et al., 2024; Liu et al., 2024c; xAI, 2023; Xiao et al., 2021; Huang et al., 2025; Hu et al., 2025; Fu et al., 2024) and ALMs (Ardila et al., 2019; Wang et al., 2021; Yang et al., 2024; Ao et al., 2024). For image modality, MMBench(Liu et al., 2024b) proposed a systematically designed benchmark to evaluate general image understanding on 20 different tasks. Focused on mathematic, MathVision(Wang et al., 2024b) collected questions from 19 mathematic competitions to evaluate VLMs complex reasoning ability. In addition to above, OCRBench(Liu et al., 2024d) supplied the evaluation on text recognition and document understanding. For video modality, MVBench(Li et al., 2024a) aggregated 11 public video benchmarks and incorporated data enhancement process to cover 20 dynamic video understanding tasks. To complement the long video understanding area, LongVideoBench(Wu et al., 2024) introduces hourly video materials to evaluate the information retrieval ability from long context. For audio modality, MMAU(Sakshi et al., 2025) provides general audio understanding assessment across speech, sounds and music domains, featuring diverse audio samples. There’s massive uni-modal benchmarks covering diverse model abilities on vision modality and audio modality separately.
2.2 Omni-Modal Benchmarks
Omni models have arisen in recent years(Comanici et al., 2025; Xu et al., 2025b; AI et al., 2025; Li et al., 2025), as the pioneer, Gemini(Comanici et al., 2025) shows a strong ability in understanding both vision and audio, while Qwen-3-Omni(Xu et al., 2025b) provides leading performance in open-source models. However, there’s less omni-modal benchmarks that can evaluate the modality combination across image, video and audio. OmniBench(Li et al., 2024b) inserted audio as a context into the image understanding task and made up an omni-modal benchmark, while the data quality needs further improvement. WorldSence(Hong et al., 2025) emphasized audio-visual data in real world scenarios with high data quality, while most audio-visual questions can be solved by audio or video alone, which can’t assess the cross-modality ability. Other datasets focus on audio (Gong et al., 2024) or video (Zhou et al., 2025) and cover limited task types.
Addressing these limitations, we propose a novel and unified benchmark, UNO-Bench, that enables comprehensive model assessment and pushing omni model to the next-level of intelligence.
@@ 3 Method
In this section, we first introduce the omni-modal dataset construction pipeline in Section.3.1. The uni-modal dataset improvement method and a general evaluation dataset compression method to improve the evaluation efficiency are introduced in Section.3.2. Finally, the multi-step open-ended questions are introduced in Section.3.3.
@@ 3.1 Omni-modal Dataset Construction
We have established a human-centric data construction pipeline (Figure.2) that efficiently empower human intelligence to produce high-quality and high-diversity dataset.
Figure 2: Dataset Construction Pipeline includes human crafted process (left side) and automated data compression (right side). First, we collect diverse and novel materials to prevent data contamination. Second, with the proposed unified ability taxonomy, human annotators including experts will craft questions, answers and record audios in real-world scenarios. Finally, with model checking, ablation study and human experts revision, we achieves high quality and diversity dataset. Regarding automated data compression, we use model performances as the sample feature, employing clustering and resampling to achieve a 90% dataset compression.
3.1.1 Model Ability Taxonomy
Through cumulative experiences on multimodal evaluation from both model-side and user-side, we summarize the capabilities of uni-modal and omni-modal into a unified model ability taxonomy. As shown in Figure.1, the omni model’s capabilities are systematically categorized into two primary dimensions: Perception and Reasoning. Detailed definitions and examples can be found in the Appendix.E.
Perception dimension structured through seven recognition types: Object Perception, Attribute Perception, Scenario Perception, Spatial Perception, Cross-Modal Conversion, Semantic Understanding. In addition, we incorporate Cross-Modal Alignment evaluation to assess information synchronization across modalities.
The perception dimension fuses visual and audio modalities. For example, Scenario Perception includes the recognition of visual scenes and the judgment of audio scenes.
Reasoning dimension extends conventional reasoning categories (General, STEM, Code) with: Spatial Reasoning (including Static Reasoning and Dynamic Reasoning), Temporal Reasoning, and Complex Reasoning (multi-conditional, multi-step problem solving).
The unified ability taxonomy combines uni-modal and omni-modal abilities which provides a comprehensive measurement that is particularly critical for advanced omni models. Based on this taxonomy, we create a diversity dataset with 44 task types distributed as shown in Figure.1.
3.1.2 Material Collection
@@
In both data quality checks and experimental results, we found that natural audio-visual synchronized video data contains a large amount of information redundancy, only a few videos require both audio and vision modality simultaneously. Therefore, we begin with carefully designed material collection. Our materials have the following three characteristics: a. Diverse Sources. The majority of our materials are real-world photos and videos collected through crowdsourcing, and another portion sourced from copyright-free website. Additionally, a small fraction comes from high-quality public datasets such as MMVU(Zhao et al., 2025), LongVideoBench(Wu et al., 2024), and VideoVista(Chen et al., 2025). b. Rich and Diverse Topics. Our materials cover a broad spectrum of subjects, including society, culture, art, life, literature, science, and so on. c. Live-Recorded Audio. Apart from background sounds and music, all dialogue is recorded by human speakers. With over 20 participants in the recording process, the audio features are rich and closely reflect the diverse vocal characteristics of the real world.
Finally, we conduct material filtering. Eliminate meaningless, illogical, and low-quality materials, and categorize the remaining materials by theme to create a material library. Additionally, label the materials with more detailed information such as subject, event, scene, and style to facilitate subsequent annotators to quickly find matching materials.
3.1.3 QA Annotation
@@
Our annotators consist of human experts and high-quality crowd-sourced users. Human experts have extensive experience in cross-modal data construction and annotation, a deeper understanding of model capabilities, and thus ensure higher professionalism and specificity in the data they construct. Most crowd-sourced users are college students with rich experience in multi-modal model interactions and diverse professional backgrounds, providing data with better authenticity and diversity.
First, annotators clarify the required image/video features based on task type definitions and filter appropriate materials from existing libraries using tags. Second, following data construction requirements, they then design prompts and corresponding answers. Third, to enhance data authenticity, all dialogue audio is recorded manually. Through this workflow, we ultimately generate complete QA pairs encompassing three modalities: visual, auditory, and textual.
Compared to conventional methods limited to human intervention only during the quality assurance phase, our pipeline integrates a "Human-in-the-Loop" approach, ensuring continuous manual involvement from the initial data sourcing to the final output. This methodology not only prevents data leakage but also more accurately simulates real-world scenarios. Furthermore, the manually curated Chinese dataset genuinely captures user requirements in a Chinese linguistic context, compensating for the shortcomings of most existing English-centric datasets.
3.1.4 Quality Inspection
@@
To ensure the data quality, we have established a multi-stage, cyclically validated quality assurance system composed of automated tools and manual review. Each question undergoes at least three rounds of independent quality inspection to maximize data quality. First, a preliminary model check is conducted to filter out cases with ambiguous questions, non-unique answers, or those that do not conform to the task type. Second, through modality ablation experiments, we remove one modality of information from the QA pair to see if the model can answer based solely on the remaining information. If the question becomes unsolvable or ambiguous after removing any one modality, it proves the cross-modality solvability of the data. Finally, manual quality inspection and revision are performed.
@@ 3.2.1 Quality Improvement
Existing public uni-modal datasets are bothered by data leakage issue(Xu et al., 2024). To verify the influence, we adopt privatization improvement on the widely used public dataset MMBench(Liu et al., 2024b). As shown in Figure.3, the performance of models have better distinguishability after dataset improvement, reflecting the true capability differences between models. Therefore, for uni-modal data, we also follow the aforementioned construction process for self-construction datasets.
Figure 3: Comparison result of privatization improvement. After improvement, the performances among models are more distinguishable.
In addition to self-constructed data, we also selected some multimodal data from public datasets to supplement in terms of capability items and data types. (Data mainly comes from AV-Odyssey(Gong et al., 2024) and WorldSense(Hong et al., 2025), accounting for 11% of the total). The specific selection logic is as follows:
Comprehensiveness: In terms of capabilities, focus on supplementing the perception part with a relatively low self-construction proportion, while also adding some reasoning questions; in terms of data types, prioritize selecting the video plus audio modality combination with a lower self-construction proportion for supplementation, followed by image plus audio.
Diversity: Supplement material types not covered in self-construction data to enhance diversity.
High Quality: Pay attention to the quality of datasets (whether uni-modal answers are reasonable and accurate).
Discriminative: Pay attention to the performance of this dataset on the model, and remove overly difficult subsets with little discrimination.